suppose two trains leave a station at the same time, traveling in opposite directions. Train B travels 20 mph faster than Train A. In 4.5 hours, the trains are 459 miles apart. Find the speed of each train.
Answer by ikleyn(52943)

.

x mph      = speed of the slower train.

(x+20) mph = speed of the faster train.


Total distance equation

    4.5x + 4.5(x+20) = 459.


Divide both sides by 4.5

    x + (x+20) = 102

      2x       = 102 - 20 = 82

       x                  = 82/2 = 41.


ANSWER.  The speed of the slower train is 41 mph.  That of the faster train is 41+20 = 61 mph.

Solved.

------------------

For simple Travel & Distance problems,  see introductory lessons
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.